OpenStudy (anonymous):

hi.Can someone please help me.I need to factor 125+64d^6 the (d) is elevated to 6 power.Thanks

OpenStudy (solomonzelman):

64d^6=(2d)^6 = [ (2d)^2 ] ^3 125 = 5^3

OpenStudy (solomonzelman):

Then use a³+b³=(a+b)(a²+ab+b²)) your "a" is (2d)^2 and your "b" is 5

OpenStudy (anonymous):

ohh,I see, I forgot that formula. Thank so much,.

OpenStudy (solomonzelman):

Anytime

OpenStudy (anonymous):

hi, can you help me with this one? (p+6)(p-2)=-7 solve

OpenStudy (solomonzelman):

1) expand the left side 2) add 16 to both sides 3) complete the square

OpenStudy (solomonzelman):

you know how to complete the square, right?

OpenStudy (anonymous):

i got p^2-2p+6p-12=-7

OpenStudy (anonymous):

why 16?

OpenStudy (anonymous):

@SolomonZelman